Find the angle 0 between the given vectors to the nearest tenth of a degree. U = – 7i +8j. V = 13i – 9;
Answer #1

Coso = Jov lullvl (-7i+ai). (138-9 il JHUJ7 (8? 12+t95? -91-72 113 1250 -163 128250 i o= cos -163 /28250 = 165.00 165.go